# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ Abstract
We conclude by outlining open challenges such as: developing femtosecond-resolved, coupled e–ph solvers, solving the full mode-to-mode Peierls–Boltzmann equation with four-particle terms, embedding correlated-electron methods (GW, dynamical mean-field theory, hybrid functionals) in e–ph workflows, implementing fully non-local, frequency-dependent screening for van-der-Waals stacks, and leveraging higher-order e–ph coupling and symmetry breaking to realise phononic thermal diodes and rectifiers. Solving these challenges will elevate electron–phonon theory from a diagnostic tool to a predictive, parameter-free platform that links symmetry, screening, and many-body effects to heat and charge transport in next-generation electronic, photonic, and thermoelectric devices.
11footnotetext: s4kazemi@uwaterloo.ca22footnotetext: gfanchin@uwo.ca11footnotetext: Corresponding author
@@ 1 Introduction
The electron-phonon (e-ph) interaction is a foundational concept in condensed matter physics, describing the interaction between charge carriers (electrons or holes) and lattice vibrations (phonons).[1, 2, 3, 4] This interaction governs a variety of phenomena in solids, including electrical resistivity[5, 6], superconductivity[7, 8], and, most pertinently for this review, thermal conductivity[9, 10, 11, 12, 13]. As modern technologies demand increasingly precise thermal management at the nanoscale, understanding the role of e-ph interaction in thermal transport, particularly in complex and low-dimensional materials, has become a critical area of research.
Thermal conductivity in solids arises from contributions of both phonons and electrons. In metals and heavily doped semiconductors, the electronic channel dominates, and electron scattering by phonons sets the thermal conductivity. By contrast, in intrinsic semiconductors and insulators, phonons are the principal heat carriers, and their lifetimes, and thus the lattice thermal conductivity, are limited by scattering with electrons or photo-excited carriers.[14, 15, 16] One can therefore decompose the total thermal resistivity into an electronic part, originating from electrons scattered by phonons, and a phononic part, originating from phonon scattering off excited carriers. The relative weight of these two contributions is dictated by carrier concentration, bonding character, and dimensionality.
@@ 1.1 From the Born–Oppenheimer approximation to its breakdown
The Born–Oppenheimer (BO) or adiabatic approximation, introduced by Born and Oppenheimer in 1927[17], assumes that ionic motion is slow enough for the electronic subsystem to remain in its instantaneous ground state. Under this premise the electronic and vibrational (phononic) Hamiltonians decouple, the e–ph interaction enters only as a weak perturbation, and charge and heat transport can be treated as two largely independent channels: phonon-mediated conduction, which dominates in insulators, and electron-mediated conduction, characteristic of metals and heavily doped semiconductors[1, 18]. Because the approximation is accurate whenever phonon frequencies are small compared with electronic energy scales, it underlies most ab initio implementations of the electron and phonon Boltzmann transport equations for bulk semiconductors and conventional metals, where e–ph coupling is moderate and nonadiabatic effects are weak.
The adiabatic picture breaks down when electronic and nuclear time scales converge, as in materials with vanishing or tunable gaps. Recently, the breakdown of the BO approximation in a new class of 2D quantum solids has been verified[19] with the Dirac crystal graphene as a progenitor of these moderately-correlated quantum systems[20, 21]. This results in the necessity of better understanding the e-ph interactions in Dirac crystals. Graphene offers a textbook example[22, 23]: its linear bands yield a low density of states at the Dirac point, and modest electrostatic gating can shift the Fermi level to values comparable with optical-phonon energies[24]. In this regime, phonon frequencies exceed the electronic momentum-relaxation rate, the electrons cannot follow the lattice adiabatically, and nonadiabatic corrections such as the experimentally observed stiffening of the G-band with doping[19] emerge. These effects can be captured by a time-dependent treatment of the phonon self-energy that keeps the real part of the e–ph self-energy, as demonstrated by Lazzeri and Mauri[25]. More generally, the BO framework inherently neglects the role of strong e–ph coupling, which can significantly alter the energy exchange dynamics between subsystems. Strong e–ph coupling or low carrier density can invalidate quasi-equilibrium models such as the two-temperature model, as shown by nonequilibrium Boltzmann simulations that include full e–e, e–ph and ph–ph scattering[26]. Accurate description of energy exchange and thermalization in such systems, therefore, demands dynamical, beyond-adiabatic frameworks that incorporate frequency-dependent screening[27] and, when necessary, higher-order electron–phonon processes.
@@ 1.2 Ab initio Frameworks for Calculating Electron–Phonon Interactions
At the core of these approaches lies Density Functional Theory (DFT) and Density Functional Perturbation Theory (DFPT)[28, 29]. DFPT gives direct access to the e–ph matrix elements $g_{mn}^{\nu}(\mathbf{k},\mathbf{q})$, the quantum amplitudes for scattering an electron from band–wave-vector state $|n\mathbf{k}\rangle$ to $|m,\mathbf{k+q}\rangle$ by absorbing or emitting a phonon $(\nu,\mathbf{q})$. Because these matrix elements must be sampled on ultra-dense ($\sim\!10^{8}$–point) Brillouin-zone grids to converge transport integrals, most modern workflows interpolate them with maximally localised Wannier functions. The e-ph coupling using Wannier functions (EPW) code[30] implements this interpolation and has become standard for both bulk (e.g. Si, Al) and layered materials such as MoS2.
Transport coefficients were first extracted by solving the Boltzmann transport equation (BTE) within the relaxation-time approximation (RTA). While the RTA offers valuable insight, it neglects momentum-conserving normal processes, drag phenomena and any dynamical coupling between the electronic and phononic distribution functions. Iterative or variational solutions of the phonon BTE[31, 32] overcome the first limitation, fully coupled e–ph BTE solvers tackle the second. A notable example is elphbolt[33, 34], which self-consistently solves the linearised electron- and phonon-BTEs under the same driving forces, thereby capturing phonon drag on electrons, electron drag on phonons and Onsager reciprocity. In doped Si, GaAs and monolayer MoS2, elphbolt reproduces time-domain thermoreflectance (TDTR) and isotope-controlled measurements to within experimental error, underscoring the importance of dynamical e–ph coupling even when the interaction itself is only moderate.
@@
In 2D Dirac crystals with low Fermi energy, and strong e–ph coupling, the standard first-order perturbative treatment (used for calculating the e-ph interaction in metals and conventional semiconductors), where phonon scattering is described by three-particle processes involving the annihilation of an e–ph pair and the creation of a new electron (commonly referred to as the $\text{EP}\!\rightarrow\!\text{E}^{\ast}$ process), can become inadequate, particularly in regimes where the Fermi energy, $E_{\mathrm{F}}$, is much smaller than the thermal energy, $\mathrm{k_{\mathrm{B}}T}$. The so-called $\text{E}\!\rightarrow\!\text{E}^{\ast}\text{P}^{\ast}$ process (electron decays into a new electron and emits a phonon) becomes non-negligible at high temperatures and low Fermi energies. This counteracts the $\text{EP}\!\rightarrow\!\text{E}^{\ast}$ scattering by reintroducing phonons into the lattice, partially canceling or strongly attenuating the net phonon depletion.[35] Therefore, at the level of three-particle interactions, the simplistic picture of phonon scattering must be corrected by including both creation and annihilation channels, particularly in undoped or lightly doped 2D Dirac materials.
In the strong-coupling limit a further correction is required: the second-order (four-particle) process $\text{EP}\!\leftrightarrow\!\text{E}^{\ast}\text{P}^{\ast}$ in which an e–ph pair is annihilated while a new pair is created. Diagrammatic analyses and kinetic-equation studies show that this channel scales with the square of the e–ph matrix element and becomes the dominant source of phonon damping at high temperature and low $E_{\mathrm{F}}$[1, 35]. Capturing it quantitatively requires both (i) a dynamic, momentum- and frequency-dependent dielectric function,[27] static Thomas–Fermi screening grossly underestimates the coupling at $q\!\simeq\!2k_{F}$, and (ii) explicit inclusion of four-particle diagrams in coupled electron– and phonon–Boltzmann solvers[35, 36]. Together these ingredients mark the crossover to a genuine strong-coupling regime in which thermal transport can no longer be described by conventional, lowest-order approximations.
@@ 1.4 Scope and Structure of This Review
By examining these case studies, we identify trends in how e-ph coupling affects thermal conductivity across different regimes, and we draw material-specific conclusions. These include the relative importance of different phonon branches, the role of carrier density and doping, and the influence of dimensionality and screening. The goal is to provide a coherent picture of when and how e–ph interactions dominate or suppress thermal transport, and what theoretical tools are best suited for capturing these effects in various classes of materials.
@@ 2.1 Constructing the Electron-Phonon Hamiltonian
contains the screened potential $\phi_{s}(q)$ whose analytic form ranges from Thomas–Fermi in weakly coupled metals to the full Lindhard response in low-density Dirac materials[4, 27, 37, 38, 39, 40, 41, 42]. In low-dimensional systems with strong coupling the four-particle channel (EP $\leftrightarrow$ EP) cannot be ignored and is treated by extending the same formalism to second order in the interaction [3, 35].

### v2

@@ Abstract
We conclude by outlining open challenges such as: developing femtosecond‑resolved, coupled e–ph solvers, solving the full mode‑to‑mode Peierls–Boltzmann equation with four‑particle terms, embedding correlated‑electron methods (GW, dynamical mean‑field theory, hybrid functionals) in e–ph workflows, implementing fully non‑local, frequency‑dependent screening for van‑der‑Waals stacks, and leveraging higher‑order e–ph coupling and symmetry breaking to realise phononic thermal diodes and rectifiers. Solving these challenges will elevate electron–phonon theory from a diagnostic tool to a predictive, parameter‑free platform that links symmetry, screening, and many‑body effects to heat and charge transport in next‑generation electronic, photonic, and thermoelectric devices.
11footnotetext: skazemi5@uwo.ca22footnotetext: gfanchin@uwo.ca11footnotetext: Corresponding author
@@ 1 Introduction
The electron-phonon (e-ph) interaction is a foundational concept in condensed matter physics, describing the interaction between charge carriers (electrons or holes) and lattice vibrations (phonons).[giustino2017electron, bai2022recent, taylor2002quantum, jishi2013feynman] This interaction governs a variety of phenomena in solids, including electrical resistivity[park2014electron, varshney2004electrical], superconductivity[einenkel2011possibility, ishihara2004interplay], and, most pertinently for this review, thermal conductivity[tong2019comprehensive, li2015electrical, liao2015ab, nika2017phonons, Lindsay2014_PRB]. As modern technologies demand increasingly precise thermal management at the nanoscale, understanding the role of e-ph interaction in thermal transport, particularly in complex and low-dimensional materials, has become a critical area of research.
Thermal conductivity in solids arises from contributions of both phonons and electrons. In metals and heavily doped semiconductors, the electronic channel dominates, and electron scattering by phonons sets the thermal conductivity. By contrast, in intrinsic semiconductors and insulators, phonons are the principal heat carriers, and their lifetimes, and thus the lattice thermal conductivity, are limited by scattering with electrons or photo‑excited carriers.[AshcroftMermin, kittel1955solid, simon2013oxford] One can therefore decompose the total thermal resistivity into an electronic part, originating from electrons scattered by phonons, and a phononic part, originating from phonon scattering off excited carriers. The relative weight of these two contributions is dictated by carrier concentration, bonding character, and dimensionality.
@@ 1.1 From the Born–Oppenheimer approximation to its breakdown
The Born–Oppenheimer (BO) or adiabatic approximation, introduced by Born and Oppenheimer in 1927[silva2018born], assumes that ionic motion is slow enough for the electronic subsystem to remain in its instantaneous ground state. Under this premise the electronic and vibrational (phononic) Hamiltonians decouple, the e–ph interaction enters only as a weak perturbation, and charge and heat transport can be treated as two largely independent channels: phonon-mediated conduction, which dominates in insulators, and electron-mediated conduction, characteristic of metals and heavily doped semiconductors[giustino2017electron, mahan2000many]. Because the approximation is accurate whenever phonon frequencies are small compared with electronic energy scales, it underlies most ab initio implementations of the electron and phonon Boltzmann transport equations for bulk semiconductors and conventional metals, where e–ph coupling is moderate and nonadiabatic effects are weak.
The adiabatic picture breaks down when electronic and nuclear time scales converge, as in materials with vanishing or tunable gaps. Recently, the breakdown of the BO approximation in a new class of 2D quantum solids has been verified[pisana2007breakdown] with the Dirac crystal graphene as a progenitor of these moderately-correlated quantum systems[keimer2017physics, calderon2020correlated]. This results in the necessity of better understanding the e-ph interactions in Dirac crystals. Graphene offers a textbook example[yan2020superconductivity, ali2016butterfly]: its linear bands yield a low density of states at the Dirac point, and modest electrostatic gating can shift the Fermi level to values comparable with optical-phonon energies[akbari2015doping]. In this regime, phonon frequencies exceed the electronic momentum-relaxation rate, the electrons cannot follow the lattice adiabatically, and nonadiabatic corrections such as the experimentally observed stiffening of the G-band with doping[pisana2007breakdown] emerge. These effects can be captured by a time-dependent treatment of the phonon self-energy that keeps the real part of the e–ph self-energy, as demonstrated by Lazzeri and Mauri[lazzeri2006nonadiabatic]. More generally, the BO framework inherently neglects the role of strong e–ph coupling, which can significantly alter the energy exchange dynamics between subsystems. Strong e–ph coupling or low carrier density can invalidate quasi-equilibrium models such as the two-temperature model, as shown by nonequilibrium Boltzmann simulations that include full e–e, e–ph and ph–ph scattering[bernardi2016first]. Accurate description of energy exchange and thermalization in such systems, therefore, demands dynamical, beyond-adiabatic frameworks that incorporate frequency-dependent screening[kazemian2023dynamic] and, when necessary, higher-order electron–phonon processes.
@@ 1.2 Ab initio Frameworks for Calculating Electron–Phonon Interactions
At the core of these approaches lies Density Functional Theory (DFT) and Density Functional Perturbation Theory (DFPT)[baroni2001phonons, giannozzi2009quantum]. DFPT gives direct access to the e–ph matrix elements $g_{mn}^{\nu}(\mathbf{k},\mathbf{q})$, the quantum amplitudes for scattering an electron from band–wave-vector state $|n\mathbf{k}\rangle$ to $|m,\mathbf{k+q}\rangle$ by absorbing or emitting a phonon $(\nu,\mathbf{q})$. Because these matrix elements must be sampled on ultra-dense ($\sim\!10^{8}$–point) Brillouin-zone grids to converge transport integrals, most modern workflows interpolate them with maximally localised Wannier functions. The e-ph coupling using Wannier functions (EPW) code[ponce2016epw] implements this interpolation and has become standard for both bulk (e.g. Si, Al) and layered materials such as MoS2.
Transport coefficients were first extracted by solving the Boltzmann transport equation (BTE) within the relaxation-time approximation (RTA). While the RTA offers valuable insight, it neglects momentum‐conserving normal processes, drag phenomena and any dynamical coupling between the electronic and phononic distribution functions. Iterative or variational solutions of the phonon BTE[fugallo2014thermal, li2014shengbte] overcome the first limitation, fully coupled e–ph BTE solvers tackle the second. A notable example is elphbolt[li2023high, protik2022elphbolt], which self-consistently solves the linearised electron‐ and phonon-BTEs under the same driving forces, thereby capturing phonon drag on electrons, electron drag on phonons and Onsager reciprocity. In doped Si, GaAs and monolayer MoS2, elphbolt reproduces time-domain thermoreflectance (TDTR) and isotope-controlled measurements to within experimental error, underscoring the importance of dynamical e–ph coupling even when the interaction itself is only moderate.
@@
1.2.1 Limitations and scope of first-principles e-ph calculations
XC-functional sensitivity.
The exchange-correlation (XC) functional influences (i) equilibrium lattice constants and internal coordinates, (ii) band structures near $E_{\mathrm{F}}$ (metals) or band edges/gaps (semiconductors/insulators), and (iii) dielectric screening. These quantities enter directly into phonon dispersions and the e-ph matrix elements $g_{mn\nu}(\mathbf{k},\mathbf{q})$, as emphasized in standard reviews and software papers on first-principles e-ph theory [Giustino2017_RMP, ponce2016epw]. In practice, generalized gradient approximation (GGA) functionals such as Perdew–Burke–Ernzerhof (PBE), provide a robust baseline for trends but tend to underbind (slightly larger lattice constants) and underestimate band gaps. The meta-GGA SCAN often yields improved structures [Sun2015_PRL_SCAN], while screened hybrids such as HSE06 [Heyd2003_JCP_HSE, Heyd2006_JCP_HSE] or GW-level quasiparticles [Onida2002_RMP_GW] provide more accurate band positions and screening. When band-gap or Fermi-surface errors dominate the uncertainty, a common mitigation is to (a) relax structures with a higher-fidelity XC (SCAN/hybrid), (b) correct the electronic spectrum via GW or a scissor shift, and (c) recompute or reweight $g_{mn\nu}$ on the corrected manifold (or assess sensitivity by finite variations of the lattice constant and gap). Throughout, we verify convergence with respect to kk/$q$-mesh density, Wannierization quality, smearing, and the acoustic sum rule [ponce2016epw, Giustino2017_RMP].  [^1] Unless noted otherwise, results surveyed here follow standard semilocal DFT/DFPT setups as reported in the cited works. Where the choice of exchange–correlation functional materially differs (e.g., meta-GGA, hybrid, GW, DFT+UU/DMFT), we indicate it explicitly.
Strongly correlated systems.
For materials with localized electrons in atomic d or f orbitals (Mott/charge-transfer insulators, Hund metals), semilocal DFT/DFPT may not capture the relevant quasiparticles or even stable phonons. In such cases, beyond-DFT frameworks are required: DFT+$U$ can correct gross localization errors [Dudarev1998_PRB_DFTU]. Density functional theory plus dynamical mean-field theory supplies temperature-dependent spectral functions and has been coupled to linear-response or frozen-phonon calculations to obtain phonons in correlated metals [Georges1996_RMP_DMFT, Kotliar2006_RMP_DMFT, Savrasov2003_PRL_DMFTphonons]; and hybrid/GW workflows improve screening and band placement for e-ph pipelines [Onida2002_RMP_GW, Heyd2003_JCP_HSE, Heyd2006_JCP_HSE]. Our analysis adopts the standard DFPT $\to$ Wannier interpolation pipeline for $g_{mn\nu}$ on dense Brillouin-zone meshes unless stated otherwise [ponce2016epw]. Where materials are known to require beyond-DFT treatments, we point this out explicitly and summarize suitable approaches (see the cross-references in the relevant sections).  [^2] A systematic treatment of strongly correlated compounds is beyond the scope of the present benchmarks but is outlined here to delimit applicability and guide future work.
1.2.2 Beyond Peierls: phonon coherences and tunneling
The semiclassical Peierls framework treats heat flow as the transport of propagating phonon wave packets (“populations”) governed by a linearized BTE. A unified quantum treatment augments this picture by including off-diagonal elements of the phonon velocity operator and the Wigner one-body density matrix, which describe wave-like tunneling (coherences) between vibrational eigenstates. Within this formalism, the thermal conductivity decomposes as
$\kappa_{\alpha\beta}\;=\;\kappa^{\mathrm{P}}_{\alpha\beta}\;+\;\kappa^{\mathrm{C}}_{\alpha\beta},$ (1)
where $\kappa^{\mathrm{P}}$ is the standard Peierls (population) term and $\kappa^{\mathrm{C}}$ arises from inter-branch coherences [Simoncelli2019_NatPhys]. The coherence channel becomes relevant when (i) phonon branches are quasi-degenerate so that $|\omega_{s}-\omega_{s^{\prime}}|$ is comparable to the linewidths $\Gamma_{s,s^{\prime}}$, and (ii) the velocity operator has sizable off-diagonal matrix elements. Under these conditions, the theory predicts a non-negligible $\kappa^{\mathrm{C}}$ contribution in complex crystals with many near-degenerate modes. Conversely, in simple crystals with well-separated branches, the coherences term is negligible and the Peierls BTE suffices.
This framework is unified in that it reduces to the Peierls result in the semiclassical limit of simple crystals and to the Allen–Feldman formula in the harmonic-glass limit, while covering intermediate regimes [Simoncelli2019_NatPhys]. As a representative example, for orthorhombic CsPbBr3 the coherences term accounts for the majority of $\kappa$ at 300 K (with $\kappa^{\mathrm{C}}\!\approx\!70\%$ of the total), whereas for silicon and diamond at 300 K $\kappa^{\mathrm{C}}$ is negligible compared to $\kappa^{\mathrm{P}}$ [Simoncelli2019_NatPhys]. In what follows, we adopt the Peierls–BTE (population) framework for all quantitative results. The coherence/tunneling term is not included unless explicitly noted. For crystals with dense, nearly-degenerate phonon spectra—where $|\omega_{s}-\omega_{s^{\prime}}|\!\sim\!\Gamma_{s,s^{\prime}}$—the unified formulation [Simoncelli2019_NatPhys] can be employed by augmenting the Peierls term with the coherences contribution $\kappa^{\mathrm{C}}_{\alpha\beta}$ constructed from off-diagonal velocity matrix elements and linewidths obtained from anharmonic (and, if relevant, e–ph) scattering.
@@ 1.2.2 Beyond Peierls: phonon coherences and tunneling
In 2D Dirac crystals with low Fermi energy, and strong e–ph coupling, the standard first-order perturbative treatment (used for calculating the e-ph interaction in metals and conventional semiconductors), where phonon scattering is described by three-particle processes involving the annihilation of an e–ph pair and the creation of a new electron (commonly referred to as the $\text{EP}\!\rightarrow\!\text{E}^{\ast}$ process), can become inadequate, particularly in regimes where the Fermi energy, $E_{\mathrm{F}}$, is much smaller than the thermal energy, $\mathrm{k_{\mathrm{B}}T}$. The so-called $\text{E}\!\rightarrow\!\text{E}^{\ast}\text{P}^{\ast}$ process (electron decays into a new electron and emits a phonon) becomes non-negligible at high temperatures and low Fermi energies. This counteracts the $\text{EP}\!\rightarrow\!\text{E}^{\ast}$ scattering by reintroducing phonons into the lattice, partially canceling or strongly attenuating the net phonon depletion.[kazemian2024influence] Therefore, at the level of three-particle interactions, the simplistic picture of phonon scattering must be corrected by including both creation and annihilation channels, particularly in undoped or lightly doped 2D Dirac materials.
In the strong-coupling limit a further correction is required: the second-order (four-particle) process $\text{EP}\!\leftrightarrow\!\text{E}^{\ast}\text{P}^{\ast}$ in which an e–ph pair is annihilated while a new pair is created. Diagrammatic analyses and kinetic-equation studies show that this channel scales with the square of the e–ph matrix element and becomes the dominant source of phonon damping at high temperature and low $E_{\mathrm{F}}$[giustino2017electron, kazemian2024influence]. Capturing it quantitatively requires both (i) a dynamic, momentum- and frequency-dependent dielectric function,[kazemian2023dynamic] static Thomas–Fermi screening grossly underestimates the coupling at $q\!\simeq\!2k_{F}$, and (ii) explicit inclusion of four-particle diagrams in coupled electron– and phonon–Boltzmann solvers[kazemian2024influence, sohier2017density]. Together these ingredients mark the crossover to a genuine strong-coupling regime in which thermal transport can no longer be described by conventional, lowest-order approximations.
@@ 1.4 Scope and Structure of This Review
Comprehensive reviews have covered first-principles e-ph theory across formalism, spectroscopy, transport, and superconductivity, and survey e-ph phenomena specifically in two-dimensional materials and device contexts [bai2022recent, Giustino2017_RMP]. Our contribution is complementary and focused: we provide a unified treatment of lattice thermal conductivity from metals to 2D Dirac crystals, emphasizing when phonon lifetimes are limited by e-ph damping, when a coupled e-ph Boltzmann treatment is essential, and when higher-order physics, four-phonon scattering, and, where warranted, coherences/unified corrections must be included. We also consolidate practical guidance (XC-functional sensitivity, beyond-DFT options) so readers can reproduce quantitative trends across these material classes rather than isolated case studies.
@@
By examining these case studies, we identify trends in how e-ph coupling affects thermal conductivity across different regimes, and we draw material-specific conclusions. These include the relative importance of different phonon branches, the role of carrier density and doping, and the influence of dimensionality and screening. Unless noted otherwise, $N(E)$ denotes the electronic density of states per energy per volume. When an explicit spin-degeneracy factor $n_{s}$ appears in an expression, $N(E)$ should be understood as per spin; otherwise it is total (including spin). The phonon density of states is written as $D_{\nu}(\omega)$ for branch $\nu$, given per angular frequency per volume (or per area in 2D). The goal is to provide a coherent picture of when and how e–ph interactions dominate or suppress thermal transport, and what theoretical tools are best suited for capturing these effects in various classes of materials.
@@ 2.1 Constructing the Electron-Phonon Hamiltonian
contains the screened potential $\phi_{s}(q)$ whose analytic form ranges from Thomas–Fermi in weakly coupled metals to the full Lindhard response in low-density Dirac materials[jishi2013feynman, kazemian2023dynamic, ziman1957effect, zhu2021dynamical, iurov2017exchange, hwang2008screening, lu2016friedel, calandra2007electron]. In low-dimensional systems with strong coupling the four-particle channel (EP $\leftrightarrow$ EP) cannot be ignored and is treated by extending the same formalism to second order in the interaction [taylor2002quantum, kazemian2024influence].